### v2

@@ Appendix B Lower bounds
Proof B.4.
By an orthogonal decomposition, $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}=\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}+\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}$. Thus, Eq. 12 can be rewritten as
$\operatorname{\mathbb{E}}\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}.$
Any algorithm that guarantees Eq. 12 also guarantees
$\operatorname{\mathbb{E}}\lVert\boldsymbol{A}^{+}\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}=\operatorname{\mathbb{E}}\lVert\boldsymbol{A}^{+}(\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star})\rVert^{2}\\ \leq\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\operatorname{\mathbb{E}}\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}.$
By Proposition B.1, the algorithm must allow access to $t\geq d/\varepsilon$ entries of $\boldsymbol{b}$.
@@ Appendix C Achieving the lower bounds: Preconditioning and initialization
Comparing the TARK error bound Theorem 1.2 to the lower bound Proposition B.1, we recognize two possible areas for improvement: the presence of the square Demmel condition number $\kappa_{\rm dem}^{2}$ in place of the dimension $d$ and the burn-in period needed to wash out the influence of the initialization $\boldsymbol{x}_{0}$. The former problem can by addressed by applying TARK to a preconditioned version of the least-squares problem
@@
The latter problem can be addressed using a careful choice of $\boldsymbol{x}_{0}\approx\boldsymbol{x}_{\star}$.
@@
Theorem C.1 (Preconditioned TARK with volume sampling).
Given a matrix $\boldsymbol{A}\in{}^{m\times d}$ of rank $r$ and a vector $\boldsymbol{b}\in{}^{m}$, consider the following algorithm:
@@
    Sample a subset of $r$ rows $\mathsf{S}\subseteq\{1,\ldots,m\}$ from the square-volume distribution [8]
    $\mathbb{P}(\mathsf{S})=\frac{\operatorname{det}(\boldsymbol{Q}(\mathsf{S},:))^{2}}{\sum_{|\mathsf{S}^{\prime}|=r}\operatorname{det}(\boldsymbol{Q}(\mathsf{S^{\prime}},:))^{2}}.$
@@
    Solve the triangular system $\boldsymbol{\widehat{x}}=\boldsymbol{R}^{+}\,\boldsymbol{\overline{y}}_{t}$, where $\boldsymbol{\overline{y}}_{t}$ is the output vector from TARK.
Then, the TARK-based solution $\boldsymbol{\widehat{x}}\in{}^{d}$ satisfies
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{\widehat{x}}\bigr{\rVert}^{2}\leq\biggl{[}1+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}r+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\cdot\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\bigr{\rVert}^{2},$
@@
Dereziński & Warmuth [8, Thm. 8] demonstrates that $\boldsymbol{y}_{0}=\boldsymbol{Q}_{\mathsf{S}}^{-1}\boldsymbol{b}_{\mathsf{S}}$ satisfies
$\operatorname{\mathbb{E}}\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{0}\right\|^{2}\leq(r+1)\cdot\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\right\|^{2}\text{ and equivalently }\operatorname{\mathbb{E}}\left\|\boldsymbol{y}_{0}-\boldsymbol{y}_{\star}\right\|^{2}\leq r\cdot\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\right\|^{2}.$
@@
$\operatorname{\mathbb{E}}\left[\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{\overline{y}}_{t}\bigr{\rVert}^{2}\,\middle|\,\boldsymbol{y}_{0}\right]$ $=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}+\operatorname{\mathbb{E}}\left[\bigl{\lVert}\boldsymbol{\overline{y}}_{t}-\boldsymbol{y}_{\star}\bigr{\rVert}^{2}\,\middle|\,\boldsymbol{y}_{0}\right]$
$\leq\biggl{[}1+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\cdot\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}\cdot\left\|\boldsymbol{y}_{0}-\boldsymbol{y}_{\star}\right\|^{2}.$
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{\overline{y}}_{t}\bigr{\rVert}^{2}\leq\biggl{[}1+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}r+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\cdot\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}.$
Since $\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{\overline{y}}_{t}\bigr{\rVert}^{2}=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{\widehat{x}}\bigr{\rVert}^{2}$ and $\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\bigr{\rVert}^{2}$, this completes the proof.
The problem of approximately solving a least-squares problem from a small number of entry evaluations of the vector $\boldsymbol{b}$ has also received recent attention in the context of active learning [6, 19]. Existing approaches achieve the guarantee $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ with high probability after accessing just $\mathcal{O}(r/\varepsilon)$ [6] or $\mathcal{O}(r\log r+r/\varepsilon)$ [27, 9] entries of $\boldsymbol{b}$. Compared to this previous work, Theorem C.1 attains nearly the optimal rate and is among the simplest and most explicit bounds for active linear regression methods.
@@ Appendix D Proofs for ridge regression
This section proves the RK-RR and TARK-RR error bounds. The analysis roughly parallels the analysis in Section 2.1. However, the proof of Theorem 3.1 requires a new strategy, since there is not a simple one-step recursion bounding $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}\right\|^{2}$ in terms of $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}$. Instead, it is necessary to use a bias–variance decomposition inspired by [7, 14].
Lemma D.1 (Multi-step expectations).
The RK-RR iteration Eq. 7 satisfies
@@
Proof D.2.
For any $t\geq 0$, rewrite the RK-RR iteration Eq. 7 as
(13) $\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}=\mu\biggl{[}\mathbf{I}-\frac{\boldsymbol{a}_{i_{t}}\boldsymbol{a}_{i_{t}}^{\top}}{\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}}\biggr{]}\bigl{(}\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\bigr{)}+\mu\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{\mu}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}-(1-\mu)\boldsymbol{x}_{\mu}.$
@@
The ridge-regularized solution $\boldsymbol{x}_{\mu}$ is characterized by $\boldsymbol{A}^{\top}\bigl{(}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\bigr{)}=\frac{1-\mu}{\mu}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}\cdot\boldsymbol{x}_{\mu}$, so the last two terms cancel. Hence, by averaging over the random indices $i_{r},\ldots,i_{s-1}$,
(14) $\operatorname{\mathbb{E}}\bigl{[}\boldsymbol{x}_{t+1}\bigr{]}=\mu\Biggl{[}\mathbf{I}-\frac{\boldsymbol{A}^{\top}\boldsymbol{A}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}}\Biggr{]}\operatorname{\mathbb{E}}\bigl{[}\boldsymbol{x}_{t}\bigr{]}\quad\text{for each }t\in\{r,\ldots,s-1\}.$
The result follows by chaining these equations together.
Lemma D.3 (Demmel condition number bound).
@@
Proof D.4.
The result follows by expanding $\boldsymbol{x}$ in $\boldsymbol{A}$’s right singular vectors.
Proof D.5 (Proof of Theorem 3.1).
To analyze the RK-RR iteration Eq. 13, introduce a bias sequence $\boldsymbol{m}_{t}$ and a variance sequence $\boldsymbol{v}_{t}$ that are recursively defined by
@@
To bound the square bias term, average over the random index $i_{t}$ and apply Lemma D.3:
@@
The analysis of the variance is more delicate. Since $\boldsymbol{v}_{t}$ follows the same recurrence as $\boldsymbol{x}_{t}$, the relation Eq. 14 from the proof of Lemma D.1 can also be applied to $\boldsymbol{v}_{t}$, yielding
@@
(15) $\operatorname{\mathbb{E}}\left\|\boldsymbol{v}_{t+1}\right\|^{2}\leq\operatorname{\mathbb{E}}\left\|\boldsymbol{v}_{t+1}+(1-\mu)\boldsymbol{x}_{\mu}\right\|^{2}\quad\text{for each }t\geq 0.$
@@
The first line uses the definition of $\boldsymbol{v}_{t+1}$, and the second and third lines the fact that $\mathbf{I}-\frac{\boldsymbol{a}_{i_{t}}^{\vphantom{\top}}\boldsymbol{a}_{i_{t}}^{\top}}{\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}}$ is an orthogonal projection matrix that is idempotent and annihilates the vector $\boldsymbol{a}_{i_{t}}$.
@@
Moreover, by averaging over the random indices $i_{0},\ldots,i_{t-1}$ and using Eq. 15,
@@
$\operatorname{\mathbb{E}}\left\|\boldsymbol{v}_{t}\right\|^{2}\leq\frac{\mu^{2}}{1-\mu^{2}}\cdot\frac{\left\|\boldsymbol{b}-\boldsymbol{A}^{\top}\boldsymbol{x}_{\mu}\right\|^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},$
which follows because $\sum_{s=1}^{\infty}\mu^{2s}=\mu^{2}/(1-\mu^{2})$. The stated result follows from the definition of $\lambda$.
Proof D.6 (Proof of Theorem 3.2).
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}=\frac{1}{(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}.$
Next analyze the terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$ for $r\leq s$ using Lemma D.1 and Theorem 3.1:
@@
$\leq\underbrace{2\mu^{r+s}(1-\kappa_{\rm dem}^{-2})^{r}\cdot\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}}_{\text{term A}}+\underbrace{\frac{2\mu^{s-r+1}}{\lambda(1+\mu)}\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\rVert^{2}}_{\text{term B}}.$
@@
$2\mu^{r+s}(1-\kappa_{\rm dem}^{-2})^{r}\cdot\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}\leq 2[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t_{\rm b}}\cdot\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}$
@@
$\operatorname{\mathbb{E}}\bigl{\lVert}\boldsymbol{\overline{x}}_{t}-\boldsymbol{x}_{\star}\bigr{\rVert}^{2}$ $=\frac{1}{(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$
$\leq 2[\mu^{2}(1-\kappa_{\rm dem}^{-2})]^{t_{\rm b}}\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}+\frac{2\mu\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\mu}}\rVert^{2}}{\lambda(1+\mu)(t-t_{\rm b})^{2}}\sum_{r,s=t_{\rm b}}^{t-1}\mu^{|s-r|}.$
@@
which completes the proof.

### v3

@@ Appendix B Lower bounds
Proof.
By an orthogonal decomposition, $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}=\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}+\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}$. Thus, (12) can be rewritten as
$\operatorname{\mathbb{E}}\,\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}.$
Any algorithm that guarantees (12) also guarantees
$\operatorname{\mathbb{E}}\,\lVert\boldsymbol{A}^{+}\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}=\operatorname{\mathbb{E}}\,\lVert\boldsymbol{A}^{+}(\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star})\rVert^{2}\leq\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\operatorname{\mathbb{E}}\,\lVert\boldsymbol{A}\widehat{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\,\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}.$
By Proposition 1, the algorithm must allow access to $t\geq d/\varepsilon$ entries of $\boldsymbol{b}$. ∎
@@ Appendix C Achieving the lower bounds: Preconditioning and initialization
We recognize two areas of improvement for TARK. First, the TARK mean square error bound in Theorem 2 depends on the square Demmel condition number $\kappa_{\rm dem}^{2}$, whereas the lower bound in Proposition 1 depends on the dimension $d$, which is always smaller. Second, the TARK error bound suggests a burn-in period is needed to wash out the influence of the initialization $\boldsymbol{x}_{0}$. The first problem can by addressed by applying TARK to a preconditioned version of the least-squares problem
@@
The second problem can be addressed using a careful choice of $\boldsymbol{x}_{0}\approx\boldsymbol{x}_{\star}$.
@@
Theorem 6 (Preconditioned TARK with volume sampling).
Given a matrix $\boldsymbol{A}\in{}^{n\times d}$ of rank $r$ and a vector $\boldsymbol{b}\in{}^{n}$, consider the following algorithm:
@@
    Sample a subset of $r$ rows $\mathsf{S}\subseteq\{1,\ldots,n\}$ from the square-volume distribution [25]
    $\mathbb{P}(\mathsf{S})=\frac{\operatorname{det}(\boldsymbol{Q}(\mathsf{S},:))^{2}}{\sum\nolimits_{|\mathsf{S}^{\prime}|=r}\operatorname{det}(\boldsymbol{Q}(\mathsf{S^{\prime}},:))^{2}}.$
@@
    Solve the triangular system $\widehat{\boldsymbol{x}}=\boldsymbol{R}^{+}\,\overline{\boldsymbol{y}}_{t}$, where $\overline{\boldsymbol{y}}_{t}$ is the output vector from TARK.
Then, the TARK-based solution $\widehat{\boldsymbol{x}}\in{}^{d}$ satisfies
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\bigr{\rVert}^{2}\leq\biggl{[}1+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}r+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\bigr{\rVert}^{2},$
@@
Dereziński & Warmuth [25, Thm. 8] demonstrates that $\boldsymbol{y}_{0}=\boldsymbol{Q}_{\mathsf{S}}^{-1}\boldsymbol{b}_{\mathsf{S}}$ satisfies
$\operatorname{\mathbb{E}}\,\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{0}\right\|^{2}\leq(r+1)\,\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\right\|^{2}\text{ and equivalently }\operatorname{\mathbb{E}}\,\left\|\boldsymbol{y}_{0}-\boldsymbol{y}_{\star}\right\|^{2}\leq r\,\left\|\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\right\|^{2}.$
@@
$\operatorname{\mathbb{E}}\,\Bigl{[}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\overline{\boldsymbol{y}}_{t}\bigr{\rVert}^{2}\,\Big{|}\,\boldsymbol{y}_{0}\Bigr{]}$ $=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}+\operatorname{\mathbb{E}}\Bigl{[}\,\bigl{\lVert}\overline{\boldsymbol{y}}_{t}-\boldsymbol{y}_{\star}\bigr{\rVert}^{2}\,\Big{|}\,\boldsymbol{y}_{0}\Bigr{]}$
$\leq\biggl{[}1+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}\,\left\|\boldsymbol{y}_{0}-\boldsymbol{y}_{\star}\right\|^{2}.$
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\overline{\boldsymbol{y}}_{t}\bigr{\rVert}^{2}\leq\biggl{[}1+\biggl{(}1-\frac{1}{r}\biggr{)}^{t_{\rm b}}r+\frac{2r-1}{t-t_{\rm b}}\biggr{]}\,\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}.$
Since $\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\overline{\boldsymbol{y}}_{t}\bigr{\rVert}^{2}=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\bigr{\rVert}^{2}$ and $\bigl{\lVert}\boldsymbol{b}-\boldsymbol{Q}\boldsymbol{y}_{\star}\bigr{\rVert}^{2}=\bigl{\lVert}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\star}\bigr{\rVert}^{2}$, this completes the proof. ∎
The problem of approximately solving a least-squares problem from a small number of entry evaluations of the vector $\boldsymbol{b}$ has also received recent attention in the context of active learning [24, 26]. Existing approaches achieve the guarantee $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ with high probability after accessing just $\mathcal{O}(r/\varepsilon)$ [24] or $\mathcal{O}(r\log r+r/\varepsilon)$ [27, 28] entries of $\boldsymbol{b}$. Compared to this previous work, Proposition 6 attains nearly the optimal rate and is among the simplest and most explicit bounds for active linear regression methods.
@@ Appendix D Proofs for ridge regression
This section proves the RK-RR and TARK-RR error bounds. The analysis roughly parallels the analysis in Section 2.1. However, the proof of Theorem 4 requires a new strategy, since there is not a simple one-step recursion bounding $\operatorname{\mathbb{E}}\,\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}\right\|^{2}$ in terms of $\operatorname{\mathbb{E}}\,\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}$. Instead, it is necessary to use a bias–variance decomposition inspired by [29, 14].
Lemma 4 (Multi-step expectations).
The RK-RR iteration (7) satisfies
@@
Proof.
For any $t\geq 0$, rewrite the RK-RR iteration (7) as
$\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}=\mu\,\biggl{[}\mathbf{I}-\frac{\boldsymbol{a}_{i_{t}}\boldsymbol{a}_{i_{t}}^{\top}}{\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}}\biggr{]}\bigl{(}\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\bigr{)}+\mu\,\frac{b_{i_{t}}-\boldsymbol{a}_{i_{t}}^{\top}\boldsymbol{x}_{\mu}}{\left\|\boldsymbol{a}_{i_{t}}\right\|^{2}}\boldsymbol{a}_{i_{t}}-(1-\mu)\boldsymbol{x}_{\mu}.$ (13)
@@
The ridge-regularized solution $\boldsymbol{x}_{\mu}$ is characterized by $\boldsymbol{A}^{\top}\bigl{(}\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\bigr{)}=\frac{1-\mu}{\mu}\left\|\boldsymbol{A}\right\|_{\rm F}^{2}\,\boldsymbol{x}_{\mu}$, so the last two terms cancel. Hence, by averaging over the random indices $i_{r},\ldots,i_{s-1}$,
$\operatorname{\mathbb{E}}\,\bigl{[}\boldsymbol{x}_{t+1}\bigr{]}=\mu\,\Biggl{[}\mathbf{I}-\frac{\boldsymbol{A}^{\top}\boldsymbol{A}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}}\Biggr{]}\operatorname{\mathbb{E}}\,\bigl{[}\boldsymbol{x}_{t}\bigr{]}\quad\text{for each }t\in\{r,\ldots,s-1\}.$ (14)
The result follows by chaining these equations together. ∎
Lemma 5 (Demmel condition number bound).
@@
Proof.
The result follows by expanding $\boldsymbol{x}$ in $\boldsymbol{A}$’s right singular vectors. ∎
Proof of Theorem 4.
To analyze the RK-RR iteration (13), introduce a bias sequence $\boldsymbol{m}_{t}$ and a variance sequence $\boldsymbol{v}_{t}$ that are recursively defined by
@@
To bound the square bias term, average over the random index $i_{t}$ and apply Lemma 5:
@@
The analysis of the variance is more delicate. Since $\boldsymbol{v}_{t}$ follows the same recurrence as $\boldsymbol{x}_{t}$, the relation (14) from the proof of Lemma 4 can also be applied to $\boldsymbol{v}_{t}$, yielding
@@
$\operatorname{\mathbb{E}}\,\left\|\boldsymbol{v}_{t+1}\right\|^{2}\leq\operatorname{\mathbb{E}}\,\left\|\,\boldsymbol{v}_{t+1}+(1-\mu)\boldsymbol{x}_{\mu}\,\right\|^{2}\quad\text{for each }t\geq 0.$ (15)
@@
The first line uses the definition of $\boldsymbol{v}_{t+1}$, and the second and third lines the fact that $\mathbf{I}-\boldsymbol{a}_{i_{t}}^{\vphantom{\top}}\boldsymbol{a}_{i_{t}}^{\top}/\lVert\boldsymbol{a}_{i_{t}}\rVert^{2}$ is an orthogonal projection matrix that is idempotent and annihilates the vector $\boldsymbol{a}_{i_{t}}$.
@@
Moreover, by averaging over the random indices $i_{0},\ldots,i_{t-1}$ and using (15),
@@
$\operatorname{\mathbb{E}}\,\left\|\boldsymbol{v}_{t}\right\|^{2}\,\leq\frac{\mu^{2}}{1-\mu^{2}}\,\frac{\left\|\boldsymbol{b}-\boldsymbol{A}^{\top}\boldsymbol{x}_{\mu}\right\|^{2}}{\left\|\boldsymbol{A}\right\|_{\rm F}^{2}},$
which follows because $\sum\nolimits_{s=1}^{\infty}\mu^{2s}=\mu^{2}/(1-\mu^{2})$. The stated result follows from the definition of $\lambda$. ∎
Proof of Theorem 5.
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\,\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\mu}\,\bigr{\rVert}^{2}=\frac{1}{(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}.$
Next analyze the terms $\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$ for $r\leq s$ using Lemmas 4 and 4:
@@
$\leq\underbrace{2\mu^{r+s}\,(1-\kappa_{\rm dem}^{-2})^{r}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}}_{\text{term A}}+\underbrace{\frac{2\mu^{s-r+1}}{\lambda\,(1+\mu)}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{\mu}\rVert^{2}}_{\text{term B}}.$
@@
$2\mu^{r+s}\,(1-\kappa_{\rm dem}^{-2})^{r}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}\leq 2\bigl{[}\mu^{2}(1-\kappa_{\rm dem}^{-2})\bigr{]}^{\,t_{\rm b}}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}$
@@
$\operatorname{\mathbb{E}}\,\bigl{\lVert}\,\overline{\boldsymbol{x}}_{t}-\boldsymbol{x}_{\star}\,\bigr{\rVert}^{2}$ $=\frac{1}{(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\operatorname{\mathbb{E}}\,\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$
$\leq 2\,\bigl{[}\mu^{2}\,(1-\kappa_{\rm dem}^{-2})\bigr{]}^{\,t_{\rm b}}\,\bigl{\lVert}\boldsymbol{x}_{0}-\boldsymbol{x}_{\mu}\bigr{\rVert}^{2}+\frac{2\mu\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\mu}}\rVert^{2}}{\lambda(1+\mu)(t-t_{\rm b})^{2}}\sum\nolimits_{r,s=t_{\rm b}}^{t-1}\mu^{\,|s-r|}.$
@@
which completes the proof. ∎